OpenStudy (anonymous):

What is the sum of the arithmetic sequence 3, 9, 15…, if there are 22 terms? a. 2,268 b. 2.028 c. 1,452 d. 1,728

OpenStudy (anonymous):

Actually, it's \[S_{n} = \frac{n}{2}(2a_{1} + [n - 1]d)\]

OpenStudy (anonymous):

Sn = 22 / 2 (2(3) + [22 - 1]6) Solve.

OpenStudy (anonymous):

Sn = 22 / 2 (2(3) + [22 - 1]6) Sn = 11 (6 + [21]6) Sn = 11 (6 + 126) Sn = 11 (132) Sn = 1452 So C. 1452 is your answer. :) And thankyou @Calcmathlete for the formula. :)

OpenStudy (anonymous):

lol. Just takes some memorization :) Btw, what grade are you in @JBrightman3 and what level math?

OpenStudy (anonymous):

I'm going into 9th grade when school starts again, but I am doing highschool honors calculus. :)

OpenStudy (anonymous):

What about you?

OpenStudy (anonymous):

That is the answer I got. Thank y`all (:

OpenStudy (anonymous):

Woah. Calculus? Can't believe that... I'm done with Algebra II and am going to Advanced Pre-calculus 10th Grade this fall.

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

No problem Tiffani. And that's cool Calcmathlete. Are you in FLVS?

OpenStudy (anonymous):

I was for Algebra II, but I'll be doing classes not math related.

OpenStudy (anonymous):

From now on anyway :)

OpenStudy (anonymous):

Ya`ll are in lower grades than me, but definitely waaay smarter! Haha (:

OpenStudy (anonymous):

lol. Is FLVS why you're at such a high level of math?

OpenStudy (anonymous):

Well last year was my first year of FLVS, I was homeschooled by my mom since 2nd grade, so I have her to thank! :)

OpenStudy (anonymous):

Ok. I know VERY few people that are at such a high level of math in 9th grade. Good Job :)